The fussy Foods deli serves a simple fruit salad with strawberries and bannanas if strawberries cost $2.30 per lb and nanners cost $.30 per lb how many lbs of both should be used to make 5 lb blend that costs $1.70 per lb.

Found 2 solutions by th3educated1, Alan3354:
Answer by th3educated1(1)

Hey there well I kinda did the problem I dont really understand the end. But I think the answer is 2 and 1/2 lbs.
Because you would want it to be even so 2 1/2 lbs of bananas and 2 1/2 lbs of strawberries would give you 5lbs.
And in the end you will be spending $6.50 cents.
I hope this helps =).

Answer by Alan3354(30993)

The fussy Foods deli serves a simple fruit salad with strawberries and bannanas if strawberries cost $2.30 per lb and nanners cost $.30 per lb how many lbs of both should be used to make 5 lb blend that costs $1.70 per lb.
----------------------------
s + n = 5 --> n = 5-s
230s + 30n = 850
Sub for n
23s + 3(5-s) = 85
23s + 15 - 3s = 85
20s = 70
s = 3.5 pounds of strawberries
n = 1.5 pounds of bananas
----------------
3.5*2.30 + 1.5*0.30
= $8.05 + $0.45
= $8.50 so it's correct